OpenStudy (anonymous):

What is 3.2 × 10−4 m/h in centimeters per second? A. 8.889 × 10−6 cm/s B. 8.889 × 10−5 cm/s C. 1.152 × 10−5 cm/s D. 1.152 × 10−7 cm/s

OpenStudy (kirbykirby):

\[ \large 3.2 × 10^{−4} \frac{m}{h}\times \frac{\frac{100~cm}{1~m}}{\frac{3600~s}{1~h}}\]